. Find the dimensions of the right circular cylinder of greatest volume which can be

inscribed in a right circular cone with radius of 5 in and height of 12 in.


1
Expert's answer
2021-05-07T10:40:06-0400

Using similar trianglr

Let h and r be the height and radius of cylinder respectively

Volume of cylinder ="\pi r^2h"


"{ 12 \over 5}={h \over 5-r}"

"h={60-12r \over 5}"


Put the value of the height into the formula for the volume of cylinder.


"V=({60-12r \over 5})\pi r^2"

"V=12\pi r^2-{{12\pi r^3} \over 5}"

differentiating V w.r.t r

"V'=24\pi r - {36\pi r^2 \over 5}"

Put V'=0

Then,

"24\pi r - {36\pi r^2 \over 5}=0"


"120 - 36r=0"

"r={120 \over 36}={10 \over 3}" in


Since we have gotten r, then

From "h={60-12r \over 5}=12-{120 \over 15}=12-8=4 in"

Greatest volume="\pi r^2h={400 \over 9}" inΒ³